# FedLog: Personalized Federated Classification with Less Communication and More Flexibility

### v1

@@ Abstract
In federated learning (FL), the common paradigm that FedAvg proposes and most algorithms follow is that clients train local models with their private data, and the model parameters are shared for central aggregation, mostly averaging. In this paradigm, the communication cost is often a challenge, as modern massive neural networks can contain millions to billions parameters. We suggest that clients do not share model parameters but local data summaries, to decrease the cost of sharing. We develop a new algorithm FedLog with Bayesian inference, which shares only sufficient statistics of local data. FedLog transmits messages as small as the last layer of the original model. We conducted comprehensive experiments to show we outperform other FL algorithms that aim at decreasing the communication cost. To provide formal privacy guarantees, we further extend FedLog with differential privacy and show the trade-off between privacy budget and accuracy.
1 Introduction
This work focuses on the design of techniques that are communication efficient and flexible for personalized federated learning (FL). The main challenge in FL is to enable communication between different clients and a server to facilitate the estimation of (personalized) aggregated models while keeping private the data in each client [1]. Hence, instead of sharing the data into a central location, clients share parameters of their local model which are then aggregated. However, this may induce large communication costs when the models are large (millions to billions of parameters) [1] and most aggregation schemes require homogeneous architectures across clients [2, 3, 4, 5, 6]. While the data never leaves the clients, the shared parameters may be thought as implicit data summaries since they indirectly capture information about the data. Given this perspective, one may ask: can we reduce communication costs by sharing small data summaries while preserving privacy and can clients utilize different model architectures?
This paper describes a new algorithm called FedLog that shares data summaries with much smaller dimensionality than model parameterization. Furthermore, clients may use different model architectures to compute those summaries. Since the parameters of those local models are never shared, they do not need to use the same architecture. Data summaries are aggregated in a central server that trains a global prediction head by Bayesian inference. More specifically, it treats the data summaries as sufficient statistics in an exponential family model and returns to the clients the canonical parameters of this exponential family model. While all clients use this same head for prediction, their locally trained models for data summarization enable them to personalize their predictions. Differential privacy (DP) is incorporated to mitigate privacy leakage via the data summaries. The paper makes the following contributions:
@@ 1 Introduction
  • FedLog: a new Federated Bayesian logistic regression algorithm, the model of which is carefully designed to provide a statistical interpretation for representation learning.
@@
  • Experiments demonstrating FedLog’s low communication cost (as small as 1.6% of the original model per round) and fast convergence under multiple situations.
@@
2 Background
2.1 Federated Learning and Personalized FL
@@
In Federated Learning (FL), there is always a collection of clients $c\in S$ who wish to collaborate, and each client holds their own data $\mathcal{D}_{c}=({\bf X}_{c},{\bf y}_{c})$ locally. We seek to train some machine learning model $\mathcal{M}_{\theta}$ with parameters $\boldsymbol{\theta}$ on these client data. Conventionally, we would centralize all the data $\mathcal{D}=\bigcup_{c\in S}\mathcal{D}_{c}$, and choose some algorithm $\mathcal{A}$ and parameter initialization $\boldsymbol{\theta}_{0}$ to learn the model:
$\mathcal{M}_{\boldsymbol{\theta}}=\mathcal{A}(\mathcal{D},\boldsymbol{\theta}_{0})$
This simple approach becomes infeasible if the clients cannot directly share their data due to privacy concerns. FL intends to tackle this problem. A trusted central server is allowed to receive and send perturbed matrices that only contain limited information about raw client data, such as model parameters. Most FL algorithms follow the following paradigm. Such algorithms iterate between local optimization and global aggregation. During local optimization at $t^{th}$ iteration, all the clients (or a proportion of the clients in some cases) start with the global consensus parameters $\theta^{t}$, and run the algorithm $\mathcal{A}$ to update the local models: $\mathcal{M}_{\boldsymbol{\theta}^{t+1}_{c}}=\mathcal{A}(\mathcal{D}_{c},\boldsymbol{\theta}^{t})$. Note the local models could have other local parameters $\tilde{\boldsymbol{\theta}}$ than $\boldsymbol{\theta}$, but is omitted in the notation for simplicity. Then, clients potentially compress their updated model parameters, and send them instead of data to the central server so that it can be aggregated into the global consensus parameters for next iteration $\boldsymbol{\theta}^{t+1}$.
@@
In this paradigm, one challenge is how to aggregate model parameters so that the resulting global consensus $\boldsymbol{\theta}^{t+1}$ is a better approximation to the centralization version. This is especially challenging when the number of local training epochs $>1$ [7], due to non-linear loss functions and predictors, thus many FL algorithms simply resort to the averaging heuristic [2, 3, 4, 5, 6]. For example, the very first FL algorithm, FedAvg [2], let the clients send all model parameters to the server without compressing, and their weighted average becomes the new global consensus.
@@
Another challenge is that clients often have different data distribution $\Pr({\bf X}),\Pr({\bf y})$ or even $\Pr({\bf y}|{\bf X})$. This is referred to as non-i.i.d. clients, and it could harm the global aggregation in terms of convergence rate and model utility [8], especially when the goal is to learn a global model. Thus, personalized FL is introduced, where a global model is not mandatory, but each client could have their own model that best fit their data distribution. See [9] for more motivation and algorithms.
2.2 Exponential Family
@@
2.3 Conjugate Priors
Suppose we wish to model the distribution of some data ${\bf x}$ by learning some parameters ${\boldsymbol{\eta}}$, and denote the likelihood induced by the model $\Pr({\bf x}|\boldsymbol{\eta})$. With Bayesian inference, $\boldsymbol{\eta}$ is often treated as random variables, and a prior distribution $\Pr(\boldsymbol{\eta})$ can be specified to incorporate prior knowledge. Then, by Bayes’ Theorem the posterior distribution is:
@@ 2.3 Conjugate Priors
The integral in Eq. 2 and thus the posterior $\Pr(\boldsymbol{\eta}|{\bf x})$ may not be tractable for arbitrary priors $\Pr(\boldsymbol{\eta})$. A convenient choice that guarantees analytical solutions is the conjugate prior. Given a likelihood $\Pr({\bf x}|\boldsymbol{\eta})$, a prior is called its conjugate prior if $\Pr(\boldsymbol{\eta})$ and $\Pr({\bf x}|\boldsymbol{\eta})$ follow the same distribution family. Specifically, if the likelihood is an exponential family, it has known conjugate priors [11].
Likelihood: $\Pr({\bf x}|\boldsymbol{\eta})=h({\bf x})\exp(\boldsymbol{\eta}^{\top}{\bf T}({\bf x})-A(\boldsymbol{\eta}))$ (3)
@@
Posterior: $\Pr(\boldsymbol{\eta}|{\bf x})=\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+{\bf T}({\bf x}),\nu+1)$ (5)
@@
2.4 Related Work
Representation learning.
The most related to our works are federated representation learning methods. These approaches separate a whole neural network into two parts, body and head. They interpret the body as a deep neural network that learns a compact feature representation from raw data, and the head as a few shallow layers that map the features to the prediction. LG-FedAvg [3] chooses to localize the body and learn local feature representations, but averages the head for global information sharing. On the contrary, FedPer [4] localizes the head but averages the body to learn a universal representation. Later, FedRep [5] also localizes the head, but instead of jointly updating the body and the head locally, they first update the local head to convergence, and then fix the head and update the body. FedProto [12] localizes the whole model, but averages feature representations by class and forces local models to learn similar representations. FedLog can also be interpreted as a representation learning algorithm, where we learn local representations by all the layers except the last one. However, unlike previous works that share model parameters and heuristically average them, we share data summaries and update the global head with Bayesian inference.
Another line of works (CCVR [13], FedPFT [14]) fit Gaussian Mixture Models (GMM) to local features extracted by a globally uniformed body, and shares the natural parameters. The server then draws virtual features from the GMM and train a global model. It is notable that these algorithms rely on pretrained FedAvg or foundation models to unify the body. Although FedLog also fits a exponential family distribution to local features, we do not need such pretrained models since our bodies are trained locally with any architecture. Also, we work with the canonical parameters (which result in smaller messages) and do not need to draw virtual samples from the learnt distribution.
Communication efficient FL.
In addition to localizing part of the original model, which decreases the message size straightforwardly, other lines of work that aim at communication efficient FL include optimization algorithms, client selection, and model compression [1]. First, it has been shown that local training epochs could impact the rounds of global communication needed [2]. Based on this observation, researchers proposed different local optimization methods to reduce the communication rounds [15, 16, 17]. On the other hand, instead of updating all clients or randomly selecting a subset of clients, researchers proposed various algorithms that accelerate the global learning process by carefully selecting clients that contribute more to the global model or are more efficient in terms of uploading time [18, 19, 20, 21]. Additionally, researchers proposed algorithms that transmit smaller messages than the whole model by reducing and compressing the parameters [22, 23, 24]. However, most such algorithms are efficient at the cost of model accuracy [25].
Bayesian FL.
Other Bayesian models have been explored to represent distributions over models and predictions in FL. The challenge is in the aggregation of the local posteriors. Various techniques have been proposed including personalized GPs [6], posterior averaging [26], online Laplace approximation [27], Thompson sampling [28], MCMC sampling [29]. There is not much in common between these techniques with our approach despite the use of Bayes theorem since prior Bayesian techniques tend to emphasize on calibration, approximating the posterior, or even different tasks.
3 Federated Bayesian Logistic Regression
In this section, we first detail our model construction and algorithm. Then, we show how to incorporate differential privacy to prevent potential privacy leakage.
3.1 FedLog
Notations.
Let $\tilde{\boldsymbol{\theta}}_{c}$ denote the parameters of a local neural network of client $c$. Let $f_{\tilde{\boldsymbol{\theta}}_{c}}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{m}$ be the function of the local neural network and $\boldsymbol{\Phi}_{c}=f_{\tilde{\boldsymbol{\theta}}_{c}}({\bf X}_{c})$ be the $m$ local features extracted from the local input. For convenience, we designate one of the features to be always $1$. Let $n_{class}$ denote the total number of classes in a classification task, and $n_{c}$ denote the size of the local dataset. Let ${\bf e}_{i}\in\mathbb{R}^{n_{class}}$ be the standard basis (i.e. one-hot vector form) of label $i\in\{1,2,\cdots,n_{class}\}$, and $\otimes:\mathbb{R}^{m}\times\mathbb{R}^{n_{class}}\rightarrow\mathbb{R}^{m*n_{class}}$ be the Kronecker product.
We assume the joint probability of each data point in $(\boldsymbol{\Phi}_{c},{\bf y}_{c})$, denoted as $\Pr(\boldsymbol{\phi},y)$, is an exponential family with canonical parameters $\boldsymbol{\eta}\in\mathbb{R}^{m*n_{class}}$ as follows.
sufficient statistics: ${\bf T}(\boldsymbol{\phi},y):=\boldsymbol{\phi}\otimes{\bf e}_{y}$ (6)
base measure: $h(\boldsymbol{\phi},y):=\frac{\exp\left(-\sum_{i=1}^{m}\boldsymbol{\phi}_{i}^{2}\right)}{\sqrt{\pi^{m}}}$ (7)
where $\boldsymbol{\phi}_{i}\in\mathbb{R}$ denotes the $i^{th}$ entry of $\boldsymbol{\phi}$. One advantage with this specific choice of $\bf T$ and $h$ is that $A$ has an explicit expression. Let $\boldsymbol{\eta}_{y}\in\mathbb{R}^{m}$ denote the $((y-1)*m)^{th}$ to $(y*m)^{th}$ entries of $\boldsymbol{\eta}$, $y\in\{1,2,\cdots,n_{class}\}$. Let $\boldsymbol{\eta}_{y,i}\in\mathbb{R}$ denote the $i^{th}$ entry of $\boldsymbol{\eta}_{y}$. Then,
@@ 3.1 FedLog
We can write out the likelihood, conditional likelihood, prior, and posterior as follows.
$\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})$ $=\frac{\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi}-\boldsymbol{\phi}^{\top}\boldsymbol{\phi})}{\sqrt{\pi^{m}}\sum_{y^{\prime}=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y^{\prime}}^{\top}\boldsymbol{\eta}_{y^{\prime}}/4)}$ (9)
$\Pr(y|\boldsymbol{\phi},\boldsymbol{\eta})$ $=\frac{\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})}{\Pr(\boldsymbol{\phi}|\boldsymbol{\eta})}=\frac{\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi})}{\sum_{y^{\prime}=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y^{\prime}}^{\top}\boldsymbol{\phi})}$ (10)
$\Pr(\boldsymbol{\eta};\boldsymbol{\chi},\nu)$ $=f(\boldsymbol{\chi},\nu)\frac{\exp(\boldsymbol{\eta}^{\top}\boldsymbol{\chi})}{(\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y^{\prime}}^{\top}\boldsymbol{\eta}_{y^{\prime}}/4))^{\nu}}$ (11)
$\Pr(\boldsymbol{\eta}|\boldsymbol{\phi},y)$ $=\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\boldsymbol{\phi}\otimes{\bf e}_{y},\nu+1)$ (12)
Unfortunately, the normalizing constant function $f(\boldsymbol{\chi},\nu)$, the marginal likelihood $\Pr(y|{\boldsymbol{\phi}})$, and the predictive posterior $\Pr(y_{*}|\boldsymbol{\phi}_{*},\boldsymbol{\phi},y)$ are still intractable. However, another advantage of our choice of ${\bf T}$ and $h$ is shown by Eq. 10: the conditional likelihood is exactly the softmax function over $\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi}$. This means we can take any deep neural network that extracts $m$ features, and append $\boldsymbol{\eta}$ as the last linear layer that maps the features to $n_{class}$ logits. Then, this composed neural network serves as a stand-alone classifier that computes the conditional likelihood $\Pr(y|{\boldsymbol{\phi}},\boldsymbol{\eta})$.

### v2

@@ Abstract
Federated representation learning (FRL) aims to learn personalized federated models with effective feature extraction from local data. FRL algorithms that share the majority of the model parameters face significant challenges with huge communication overhead. This overhead stems from the millions of neural network parameters and slow aggregation progress of the averaging heuristic. To reduce the overhead, we propose to share sufficient data summaries instead of raw model parameters. The data summaries encode minimal sufficient statistics of an exponential family, and Bayesian inference is utilized for global aggregation. It helps to reduce message sizes and communication frequency. To further ensure formal privacy guarantee, we extend it with differential privacy framework. Empirical results demonstrate high learning accuracy with low communication overhead of our method.
Introduction
Representation learning plays a crucial role in machine learning by effectively extracting features from raw data to facilitate downstream prediction in various fields (Liu, Shen, and Yang 2024). Combined with federated learning (FL), federated representation learning (FRL) aims to learn personalized federated models and diminish the impact of heterogeneous clients. FRL methods separate a whole neural network into two parts, body and head (Liang et al. 2020; Collins et al. 2021; Arivazhagan et al. 2019). The body is a deep network that learns a compact feature representation from the raw data. The head is a shallow network with few layers that make predictions in the representation space. Personalization is achieved by localizing either the body or the head. Parameters of the rest of the model are shared with the server for global aggregation with the FedAvg heuristic (McMahan et al. 2017).
FRL utilizing body parameter sharing faces two significant challenges. 1) Heavy communication overhead. Sharing body parameters induces huge communication cost per round due to millions of parameters of the deep learning model (Wen et al. 2023) since the body contains the majority of the parameters. Furthermore, the averaging heuristic slows down the convergence with heterogeneous clients, resulting in more aggregation rounds. 2) Rigid model architecture. Sharing body parameters requires the same body architecture among clients (McMahan et al. 2017; Arivazhagan et al. 2019; Collins et al. 2021; Zhang et al. 2024; Oh, Kim, and Yun 2021; Li, Li, and Varshney 2021). However, clients usually possess different amounts of computing resources. Thus, one body architecture may not be suitable for all devices. In this case, clients with limited resources cannot effectively participate. Even though some recent works consider sharing head parameters to reduce communication overhead, their communication efficiency is still limited due to slow averaging (Karimireddy et al. 2020).
To tackle the above challenges, our motivation is that a succinct data sharing protocol should minimize the bandwidth usage and does not depend on specific model architecture. Note that sharing model parameters in original FRL can be interpreted as sharing implicit client data summaries, since local model parameters capture the information from the input which can recover the raw data (Mothukuri et al. 2021). However, there is no guarantee that the information reflected by these parameters is sufficient or necessary to infer the global model. Instead, we can consider sharing concise, but sufficient client data summaries.
Sharing sufficient data summaries (also known as sufficient statistics (Jordan 2009a)) of clients offers two benefits. 1) sufficient statistics maintain small data sizes which reduce the bandwidth usage. 2) sufficient statistics are model independent, which allows heterogenous model deployment. Based on this idea, we propose Federated Bayesian Logistic Regression (FedLog), a new FRL strategy. We consider the case that the body of the model is localized and the head is updated by the server. FedLog acquires sufficient statistics from the client data encoded by the body with an exponential family distribution. The sufficient statistics from each client are then sent to the server to determine the optimal head parameters by maximizing the posterior with Bayesian inference. The theoretical property of sufficient statistics ensures that sufficient information is captured with fixed size to infer global model parameters. Also, due to the Bayesian inference step, the number of communication rounds is reduced thus improving communication efficiency. Note that since the model body is not shared and summation of sufficient statistics is non-invertible, FedLog also avoids potential privacy attacks through weight manipulation, GAN-based reconstruction, or large model memorization effects (Boenisch et al. 2021; Mothukuri et al. 2021). However, to ensure a formal privacy guarantee, we further incorporate differential privacy framework to mitigate privacy leakage.
In summary, the paper makes the following contributions:
@@ Introduction
  • FedLog: a new FRL algorithm, the model of which is carefully designed to provide a statistical interpretation for representation learning.
@@
  • Experiments demonstrating FedLog’s low communication cost (as small as $0.09\%$ of FedAvg) and fast convergence under multiple scenarios.
@@
Related Works
Federated Representation Learning
@@
In FL, there is a collection of clients $c\in S$ wishing to collaborate. Each client holds their own data $\mathcal{D}_{c}=({\bf X}_{c},{\bf y}_{c})$ locally. We seek to train some ML model with parameters $\boldsymbol{\theta}$ on these client data. Conventionally, we would centralize all the data $\mathcal{D}=\bigcup_{c\in S}\mathcal{D}_{c}$ and learn the model with $\mathcal{D}$. This approach becomes infeasible if the clients cannot share their data due to privacy concerns. FL intends to tackle this problem. A trusted central server is allowed to receive and send perturbed matrices that only contain limited information about raw data, such as model parameters.
@@
One challenge of FL is how to aggregate model parameters learnt locally so that the resulting global consensus $\boldsymbol{\theta}^{t+1}$ is a better approximation to the centralization version than the last round $\boldsymbol{\theta}^{t}$ This is especially challenging when the number of local training epochs $>1$ (Karimireddy et al. 2020), due to non-linear loss functions and predictors. Thus, many FL algorithms simply resort to the averaging heuristic (McMahan et al. 2017; Liang et al. 2020; Arivazhagan et al. 2019; Collins et al. 2021; Achituve et al. 2021).
@@
In FL, clients often have different data distribution $\Pr_{c}({\bf X}),\Pr_{c}({\bf y})$ or even $\Pr_{c}({\bf y}|{\bf X})$. This is referred to as heterogeneous or non-i.i.d. clients. The averaging heuristic can drastically harm the global aggregation in terms of convergence rate and model utility with non-i.i.d. clients (Li et al. 2019). Thus, personalized FL (PFL) is introduced, where a global model is not mandatory, but each client could have their own model that best fits their data distribution (Tan et al. 2022a). Nevertheless, most algorithms still utilize the averaging heuristic for aggregation.
One line of FRL works approach PFL by localizing either the body (LG-FedAvg (Liang et al. 2020)) or the head (FedPer (Arivazhagan et al. 2019), FedRep (Collins et al. 2021)) of the client models, and shares the rest with the server for averaging. FedProto (Tan et al. 2022b) localizes the whole model, but averages feature representations by class and forces local models to learn similar representations. FedLog can also be interpreted as a representation learning algorithm, where we learn local representations by all the layers except the last one. However, unlike previous works that share model parameters and heuristically average them, we share sufficient statistics and update the global head with Bayesian inference.
Another line of works (CCVR (Luo et al. 2021), FedPFT (Beitollahi et al. 2024)) fit Gaussian Mixture Models (GMM) to local features extracted by a globally uniformed body, and shares the natural parameters. The server then draws virtual features from the GMM and trains a global model. It is notable that these algorithms rely on pretrained FedAvg or foundation models to unify the body. Although FedLog also fits an exponential family distribution to local features, we do not need such pretrained models since our bodies are trained locally with any architecture. Also, we work with the canonical parameters and do not need to draw virtual samples from the learnt distribution.
Other related works include communication efficient FL and Bayesian FL. See Appendix A for more details.
Method
Exponential Family
We start by introducing the definition of exponential family.
Definition 1.
@@
where $h({\bf x}):\mathbb{R}^{p}\rightarrow\mathbb{R}_{\geq 0}$, ${\bf T}({\bf x}):\mathbb{R}^{p}\rightarrow\mathbb{R}^{d}$, $A(\boldsymbol{\eta}):\mathbb{R}^{d}\rightarrow\mathbb{R}$ are known functions.
@@
Many well-known distributions are included in the exponential family, such as Gaussian, binomial, Poisson, and Bernoulli distribution. We can always transform a distribution represented by natural parameters into its canonical form defined above. For example, a binomial distribution has the following p.d.f.:
$\Pr(x|p)={n\choose x}p^{x}(1-p)^{(n-x)},x\in\{0,1,\cdots,n\}$
It can be rewritten as:
$\Pr(x|\eta)={n\choose x}\exp({\eta}x-n\ln({1+e^{\eta}})),\eta=\ln\frac{p}{1-p}$
Exponential family has a few desirable properties, which makes it a perfect candidate for information sharing in PFL. 1) The sufficient statistic ${\bf T}({\bf x})$ captures all the information of ${\bf x}$ that can be used to infer $\boldsymbol{\eta}$ (i.e., ${\bf x}{\perp\!\!\!\perp}\boldsymbol{\eta}|{\bf T(x)}$, conditional independence) (Jordan 2009a). 2) If ${\bf x}_{1},{\bf x}_{2},\cdots,{\bf x}_{n}$ are i.i.d. samples from $\Pr({\bf x}|{\boldsymbol{\eta}})$, the summation of sufficient statistics $\sum_{i=1}^{n}{\bf T}({\bf x}_{i})$ is a complete statistic for $\boldsymbol{\eta}$. It contains only information about $\boldsymbol{\eta}$, without ancillary information (Casella and Berger 2015). 3) It is the only parametric distribution family with sufficient statistics of fixed size that does not grow with the sample size (Koopman 1936). 4) It has known conjugate priors (Jordan 2009b), a crucial property for Bayesian inference as we will elaborate later. Formal definitions of these properties are in Appendix B.
Model Construction
We detail our model construction and algorithm with the following notation.
Notation.
Let $\tilde{\boldsymbol{\theta}}_{c}$ denote the parameters of a local neural network of client $c$. This network serves as a local body. Let $f_{\tilde{\boldsymbol{\theta}}_{c}}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{m}$ be the function of the local neural network and $\boldsymbol{\Phi}_{c}=f_{\tilde{\boldsymbol{\theta}}_{c}}({\bf X}_{c})$ be the $m$ local features extracted from the local input. For convenience, we designate the first feature to be always $1$. Let $n_{class}$ denote the total number of classes in a classification task, and $n_{c}$ denote the size of the local dataset. Let ${\bf e}_{i}\in\mathbb{R}^{n_{class}}$ be the standard basis (i.e. one-hot vector form) of label $i\in\{1,2,\cdots,n_{class}\}$, and $\otimes:\mathbb{R}^{m}\times\mathbb{R}^{n_{class}}\rightarrow\mathbb{R}^{m*n_{class}}$ be the Kronecker product.
We assume the joint probability of each data point in $(\boldsymbol{\Phi}_{c},{\bf y}_{c})$, denoted as $\Pr(\boldsymbol{\phi},y)$, is an exponential family with canonical parameters $\boldsymbol{\eta}\in\mathbb{R}^{m*n_{class}}$. We design the sufficient statistics ${\bf T}(\boldsymbol{\phi},y)$ and the base measure $h(\boldsymbol{\phi},y)$ to have the following form.
${\bf T}(\boldsymbol{\phi},y)$ $:=\boldsymbol{\phi}\otimes{\bf e}_{y}$ (2)
$h(\boldsymbol{\phi},y)$ $:=\frac{\exp\left(-\sum_{i=1}^{m}\boldsymbol{\phi}_{i}^{2}\right)}{\sqrt{\pi^{m}}}$ (3)
where $\boldsymbol{\phi}_{i}\in\mathbb{R}$ denotes the $i^{th}$ entry of $\boldsymbol{\phi}$. One advantage with this specific choice of $\bf T$ and $h$ lies in the resulting conditional likelihood $\Pr(y|\boldsymbol{\phi},\boldsymbol{\eta})$. Let $\boldsymbol{\eta}_{y}\in\mathbb{R}^{m}$ denote the $((y-1)*m)^{th}$ to $(y*m)^{th}$ entries of $\boldsymbol{\eta}$, $y\in\{1,2,\cdots,n_{class}\}$. Then,
$\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})$ $=\frac{\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi}-\boldsymbol{\phi}^{\top}\boldsymbol{\phi}-A({\boldsymbol{\eta}}))}{\sqrt{\pi^{m}}}$ (4)
$\Pr(y|\boldsymbol{\phi},\boldsymbol{\eta})$ $=\frac{\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})}{\Pr(\boldsymbol{\phi}|\boldsymbol{\eta})}=\frac{\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})}{\sum_{y^{\prime}=1}^{n_{class}}\Pr(\boldsymbol{\phi},y^{\prime}|\boldsymbol{\eta})}$ (5)
$=\frac{\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi})}{\sum_{y^{\prime}=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y^{\prime}}^{\top}\boldsymbol{\phi})}$ (6)
Eq. 6 is exactly the softmax function over $\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi}$. This means we can take any deep neural network that extracts $m$ features, and append $\boldsymbol{\eta}$ as the last linear layer that maps the features to $n_{class}$ logits. Then, this composed neural network serves as a stand-alone classifier that computes the conditional likelihood $\Pr(y|{\boldsymbol{\phi}},\boldsymbol{\eta})$. However, Eq. 6 cannot be directly maximized at the server, since it requires knowledge of uncompressed representation-label pairs. Instead, we utilize Bayesian inference to optimize $\boldsymbol{\eta}$.
With Bayesian inference, $\boldsymbol{\eta}$ is treated as a random variable. A prior distribution $\Pr(\boldsymbol{\eta})$ can be specified to incorporate prior knowledge. Then, by Bayes’ Theorem, the posterior distribution is:
$\Pr(\boldsymbol{\eta}|\boldsymbol{\phi},y)$ $=\frac{\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})\Pr(\boldsymbol{\eta})}{\Pr(\boldsymbol{\phi},y)}$ (7)
@@
The integral in Eq. 8 and thus the posterior $\Pr(\boldsymbol{\eta}|\boldsymbol{\phi},y)$ may not be tractable for arbitrary priors $\Pr(\boldsymbol{\eta})$. A convenient choice that guarantees analytical solutions is the conjugate prior. Given a likelihood $\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})$, a prior is called its conjugate prior if $\Pr(\boldsymbol{\eta})$ and $\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})$ follow the same distribution family. Specifically, if the likelihood is an exponential family, it has known conjugate priors (Jordan 2009b).
@@
Posterior: $\Pr(\boldsymbol{\eta}|\boldsymbol{\phi},y)=\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+{\bf T}(\boldsymbol{\phi},y),\nu+1)$ (10)
@@
Another advantage of our model design is that $A$ has an explicit expression. Let $\boldsymbol{\eta}_{y,i}\in\mathbb{R}$ denote the $i^{th}$ entry of $\boldsymbol{\eta}_{y}$.
@@
Due to this analytical solution, we can directly optimize Eq. 10 without further approximations.
FedLog
Based on the above model, we propose our new algorithm FedLog, summarized in Algo. 1. At the beginning, the server initializes $\boldsymbol{\eta}$ (the global head) randomly. The clients initialize $\tilde{\boldsymbol{\theta}}_{c}$ (the local bodies) either completely randomly, or with the same random seed sent by the server to unify the initialization. Note $\tilde{\boldsymbol{\theta}}_{c}$ is not part of our exponential family assumption, thus we do not require them to have the same shape or architecture amongst different clients.
$\begin{split}\mathcal{L}_{c}&=-\sum_{i=1}^{n_{c}}\ln\Pr({\bf y}_{c,i}|\boldsymbol{\Phi}_{c,i},\boldsymbol{\eta})\\ &=-\sum_{i=1}^{n_{c}}\ln\frac{\exp(\boldsymbol{\eta}_{y_{c,i}}^{\top}\boldsymbol{\Phi}_{c,i})}{\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\Phi}_{c,i})}\end{split}$ (12)
where $(\boldsymbol{\Phi}_{c,i},{\bf y}_{c,i})\in\mathbb{R}^{m}\times\mathbb{R}$ is the $i^{th}$ data point of $\boldsymbol{\Phi}_{c},{\bf y}_{c}$. This is exactly the cross entropy loss widely used in deep learning for classification tasks. Then, clients compute sufficient statistics of their local data $\sum_{i=1}^{n_{c}}{\bf T}(\boldsymbol{\Phi}_{c,i},{\bf y}_{c,i})=\sum_{i=1}^{n_{c}}\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}}$. They then send the summations and $n_{c}$ to the server for global head learning. As we have discussed with the introduction of exponential family, the sufficient statistics contain all information in the representations that could be used to infer $\boldsymbol{\eta}$ in our model. The server only needs to know the summation of all the sufficient statistics, since
$\begin{split}&\Pr(\boldsymbol{\eta}|\boldsymbol{\Phi}_{c_{1}},{\bf y}_{c_{1}},\cdots,\boldsymbol{\Phi}_{c_{k}},{\bf y}_{c_{k}})\\ &=\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\sum_{c\in S}\sum_{i=1}^{n_{c}}\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}},\nu+\sum_{c\in S}n_{c})\end{split}$ (13)
$\begin{split}\boldsymbol{\eta}&=\operatorname*{arg\,max}_{\boldsymbol{\eta}}\ln\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\boldsymbol{\Phi},\nu+n)\\ &=\operatorname*{arg\,max}_{\boldsymbol{\eta}}\ln\frac{\exp\left(\boldsymbol{\eta}^{\top}(\boldsymbol{\chi}+\boldsymbol{\Phi})\right)}{(\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\eta}_{y}/4))^{(\nu+n)}}\end{split}$ (14)